OpenStudy (anonymous):

If 15m = 63, find the value of 4m. A. 16.8 B. 0.95 C. 4.2 D. 7.3 If 15m = 63, find the value of 4m. A. 16.8 B. 0.95 C. 4.2 D. 7.3 @Mathematics

OpenStudy (anonymous):

15m=63 m=4.2 4*m=4*4.2=16.8 so A